### v1

# First Study of the Nuclear Response to Fast Hadrons via Angular Correlations between Pions and Slow Protons in Electron–Nucleus Scattering

Abstract
We report on the first measurement of angular correlations between high‑energy pions and slow protons in electron–nucleus ($eA$) scattering, providing a new probe of how a nucleus responds to a fast‑moving quark. The experiment employed the CLAS detector with a 5-GeV electron beam incident on deuterium, carbon, iron, and lead targets. For heavier nuclei, the pion-proton correlation function is more spread-out in azimuth than for lighter ones, and this effect is more pronounced in the $\pi p$ channel than in earlier $\pi\pi$ studies. The proton‑to‑pion yield ratio likewise rises with nuclear mass, although the increase appears to saturate for the heaviest targets. These trends are qualitatively reproduced by state‑of‑the‑art $eA$ event generators, including BeAGLE, eHIJING, and GiBUU, indicating that current descriptions of target fragmentation rest on sound theoretical footing. At the same time, the precision of our data exposes model–dependent discrepancies, delineating a clear path for future improvements in the treatment of cold-nuclear matter effects in $eA$ scattering.
1 Introduction
The production and propagation of hadrons through nuclei remain central topics in quantum chromodynamics (QCD). Key questions include: “How does the process of hadronization manifest within a nucleus?”, and “How are different hadrons produced in a single scattering event correlated with each other?” LRP2023. Addressing these questions calls for detailed studies of diverse reaction channels, encompassing multiple hadron flavor combinations in the final state and spanning a broad kinematic range, to probe the different timescales of hadronization.
Despite extensive theoretical and phenomenological work on nucleon production in nuclear deep-inelastic scattering (DIS) E665:1994aiu, Degtyarenko:1997, BEBCWA59:1989ayi, CiofidegliAtti:2004pv, Palli:2009it, Larionov:2018igy, PhysRevD.100.073010, Chang:2022hkt, RoblesGajardo:2022efe, no dedicated study has correlated the production of low-energy protons with the kinematics of leading hadrons, whose spectra are known to be modified by cold nuclear matter effects as observed by the HERMES HERMES:2000ytc, HERMES:2003icw, Airapetian:2007vu, Airapetian:2009jy, Airapetian:2011jp and CLAS CLAS:2021jhm, CLAS:2022oux, CLAS:2012tlh, kbhz-h4jv collaborations. The present work addresses this gap by employing pion-proton correlations over a wide azimuthal and rapidity range, thereby linking studies of current and nuclear fragmentation in DIS off deuterium and heavy nuclei.
In this study, we examine electron–nucleus ($eA$) DIS events that contain a leading $\pi^{+}$ accompanied by a secondary proton, as illustrated in Fig. 1. Correlating these two hadrons provides new insight into how the nucleus responds to a fast quark or hadron and how the hadronization process is modified within the nucleus. Most of the observed protons come from fragmentation of the nuclear remnant. Because the proton’s rest energy is already present in the initial state, a smaller amount of energy from the hadronic cascade is necessary to knock a proton out of a nucleus than produce a secondary pion with the same momentum. This allows the current measurement to probe lower energy scales of the hadronization process and transport properties of the nucleus than in existing dipion measurements PhysRevC.111.035201.
Figure 1: Illustration of nuclear deep‑inelastic scattering featuring a leading charged pion and secondary protons in the final state.
Protons produced through nuclear fragmentation, historically referred to as “grey” protons, were investigated in earlier scattering experiments with both hadronic and leptonic beams E910:1999ozb, E665:1994aiu, Degtyarenko:1997. Measurements of slow protons, together with model calculations, served to estimate the collision centrality, expressed as the mean number of projectile–nucleon interactions. Lepton measurements offer a cleaner test of intra-nuclear hadron-cascade models; in hadron collisions, multiple projectile–nucleon interactions complicate particle‐production mechanisms. Similar studies were carried out with neutrino DIS on nuclear targets BEBCWA59:1989ayi, although with much more limited precision.
Ciofi degli Atti et al. CiofidegliAtti:2004pv, Palli:2009it proposed that the dependence of slow-proton production over a wide kinematic range can serve to discriminate among hadronization models, and their calculation reproduced well the data from Ref. E665:1994aiu. Larionov et al. Larionov:2018igy showed that slow-neutrons provide similar insight into intranuclear cascades.
The study of slow protons (or more generally, slow nucleons) will also play an important role in the future Electron-Ion Collider Accardi:2012qut, AbdulKhalek:2021gbh, since slow nucleons will be measured with dedicated detectors, to enable various studies such as spectator tagging in electron-deuteron collisions, centrality estimations, and hadronization studies in $eA$ collisions AbdulKhalek:2021gbh, Chang:2022hkt, RoblesGajardo:2022efe.
2 Experimental setup
[…]
The CLAS CLAS:2003umf detector was based on a six-fold symmetric toroidal magnet, which defined six sectors instrumented with drift chambers (DC), time-of-flight scintillation counters (TOF), Cherenkov counters (CC), and an electromagnetic calorimeter (EC). Its polar angular acceptance ranged from 8 to 140 for the drift chambers and TOF, and from 8 to 45 for the Cherenkov counters and calorimeter. The drift chambers functioned in a toroidal magnetic field up to 2 Tm, where the magnet’s polarity made negatively charged particles bend toward the beam pipe. The resulting resolutions for charged particles were $\sigma_{p}/p=0.5\%$ for the momentum, 2 mrad for the polar angle, and 4 mrad for the azimuthal angle. The data were selected with a trigger that required at least one electron candidate with momentum $p>500$ MeV/c/c.
3 Event selection and observables
Following Refs. CLAS:2012tlh, CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201, electrons were identified by matching negatively charged tracks measured in the DC with hits in the TOF and EC. Background from $\pi^{-}$ was suppressed to the $1\%$ level using the CC and the EC. Charged pions and protons were identified with TOF hits consistent with those types of hadrons with momentum determined in the DCs. The selection for $\pi^{+}$ with momentum above 2.7 GeV/c/c was further refined with the CC to suppress proton background. Fiducial cuts on momentum and angles were used in order to avoid regions with steeply varying acceptance or low resolution. The protons’ energies were corrected for energy loss in the target using the same procedure as was used in Ref. CLAS:2022oux.
Similar to Refs. CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201, we selected events with $Q^{2}>1$ $2W>2 GeV/c2c^{2}$ and $2.3<\nu<4.2$ GeV. Here, $Q^{2}$ is the square of the four-momentum transfer from the electron to the struck nucleon,
$W=\allowbreak\sqrt{2m_{p}\nu+m_{p}^{2}-Q^{2}}$ (1)
is the magnitude of the four-momentum of the virtual photon+struck nucleon system, where $m_{p}$ is the mass of a proton, $\nu=E-E^{\prime}$ is the energy transfer, and $E$ and $E^{\prime}$ are the beam- and scattered-electron energies.
We selected events with a “leading” pion, defined as having fractional energy $z_{1}=E_{h}/\nu>0.5$ HERMES:2005mar, CLAS:2022asf, PhysRevC.111.035201, where $E_{h}$ is the energy of the pion. Following Refs. CLAS:2022oux, Hen:2014nza, we required the proton to have momentum greater than 350 MeV/c/c so that it could be well-reconstructed. The maximum proton momentum in our accepted $\pi p$ event sample was about 2.7 GeV/c/c. In order to ensure that the azimuthal angles of both particles are well defined, we further required that both the proton and the leading pion have transverse momenta (defined with respect to the momentum-transfer vector $\vec{q}$) greater than 70 MeV/c/c.
We selected events within the range $0.0<\Delta Y<3.0$, where $\Delta Y\equiv Y_{\pi}-Y_{p}$ is the rapidity separation between the pion and the proton. Here, $Y_{\pi}$ and $Y_{p}$ are the rapidities of the pion and the proton, respectively, each given by
$Y_{h}=\frac{1}{2}\log\frac{E_{h}+p_{z,h}}{E_{h}-p_{z,h}},$ (2)
where $E_{h}$ is the hadron’s energy and $p_{z,h}$ is its momentum in the momentum-transfer direction.
We selected particles arising from scattering from either the deuterium or nuclear targets by using the longitudinal vertex position defined by intersecting their trajectories with the beamline. The resulting vertex resolution ensured negligible ambiguity in the target tagging CLAS:2021jhm.
We used the electron, the leading pion, and the proton variables to measure the 2-dimensional correlation function, defined as
$C(\Delta\phi,\Delta Y)=C_{0}\times\frac{1}{N_{e^{\prime}\pi}}\left(\frac{d^{2}N_{e^{\prime}\pi p}}{d\Delta\phi\,d\Delta Y}\right),$ (3)
where $N_{e^{\prime}\pi}$ is the number of events with a scattered electron and a leading $\pi^{+}$, and $\Delta\phi\equiv\phi_{\pi}-\phi_{p}$ is the azimuthal separation between the pion and proton, as defined around the momentum-transfer axis. We determined the expression in parentheses by taking $N_{e^{\prime}\pi p}$, the number of events with a scattered electron, a leading $\pi^{+}$ and a proton in each bin in $|\Delta\phi|$ and $\Delta Y$, and dividing that by the bin widths in both variables ($\pi/8$ and 0.5, respectively), and by a factor of two, since each $|\Delta\phi|$ bin covers two $\Delta\phi$ bins. These correlation functions are corrected for pair-acceptance effects using Monte Carlo (MC) simulations. Following Ref. PhysRevC.111.035201, we also used a data-driven method to correct for contamination of the deuterium sample due to scattering off the aluminum endcap walls of the liquid deuterium cells. Details for both of these corrections can be found in the Supplementary Materials supplementary.
The normalization factor $C_{0}$ was chosen such that the integral of $C(\Delta\phi,\Delta Y)$ over $\Delta\phi$ and $\Delta Y$ equals unity for deuterium. We used the same value of $C_{0}$ for all targets. Explicitly, the formula we used was
$C_{0}=\frac{N^{D}_{e^{\prime}\pi}}{N^{D,\rm tot}_{e^{\prime}\pi p}}$ (4)
where $N^{D}_{e^{\prime}\pi}$ is the number of events from the deuterium dataset with the scattered electron and leading pion, and $N^{D,\rm tot}_{e^{\prime}\pi p}$ is the total number of events in the deuterium dataset with the scattered electron, leading pion, and proton, for all kinematic bins.
From here, we determined the nuclear-to-deuterium ratio of the correlation functions, $R=C_{A}/C_{D}$, which can be used to compare the correlation functions of the nuclear targets with those of deuterium.
We determined from these correlation functions the azimuthal width $\sigma$, defined by PhysRevC.111.035201
$\sigma=\sqrt{\frac{\sum\limits_{i\in\rm bins}C_{i}\left(\Delta\phi_{i}-\pi\right)^{2}}{\sum\limits_{i\in\rm bins}C_{i}}},$ (5)
as well as the azimuthal broadening PhysRevC.111.035201 (using carbon as our reference target [^2] We chose carbon instead of deuterium (as was done in Ref. PhysRevC.111.035201) for the reference target here, since the results for the nuclear targets are more similar to one another than they are to deuterium.),
$b=\pm\sqrt{\left|\sigma_{A}^{2}-\sigma_{C}^{2}\right|},$ (6)
where the sign of the $\pm$ is that of the expression inside the absolute value. This definition of $b$ as a quadrature difference was a natural choice given that the standard deviation of the convolution of two distributions is the quadrature sum of their standard deviations.
4 Uncertainties
We performed studies on various possible sources of systematic uncertainties using data and simulation studies. For the simulation studies we used the PYTHIA6 event generator and the GSIM package GSIM, which is based on GEANT3 Brun:1994aa [^3] It should be noted that Geant4 GEANT4:2002zbu had not been released yet when GSIM was being developed, and that GEANT3 was still the standard at the time that the data were taken., to simulate the response of the CLAS detector and dual-target setup Hakobyan:2008kua. The simulation was tuned to provide a reasonable description of the data. The possible sources of systematic uncertainties studied include acceptance effects, event selection, and particle misidentification. For the azimuthal RMS widths and broadenings, there was a small systematic uncertainty due to the finite bin width used to calculate them. Other sources of systematic uncertainty, such as cross-contamination between bins, beam luminosity, trigger efficiency, Coulomb effects, and time-dependent effects, were found to be negligible. We show the contributions of various types of systematic uncertainties for the correlation functions and the ratio $R$ in Table 4 and for the RMS widths and broadenings in Table 4. Each of the listed uncertainties are “point-to-point”, affecting each bin by a different amount (although there may be some correlation between the bins).
The total relative systematic uncertainty ranged from 15-41% for $C(\Delta\phi,\Delta Y)$, 6-21% for $R$, and 3-6% for $\sigma$. For $b$, the systematic uncertainty was about 4% for almost all bins [^4] The one exception is the highest $\Delta Y$ bin for Fe, where both the statistical and systematic uncertainties are much larger than usual. The reason for this is that the measured $\sigma$ values for C and Fe are very close together, much less than one statistical error bar. Therefore, the uncertainties on their quadrature difference are very large.. Details of these studies on systematic uncertainties can be found in the Supplementary Materials supplementary.
Source $\Delta C/C$ (D) $\Delta C/C$ (A) corr. D vs A $\Delta R/R$
Statistics 1-44% 2-41% N 2-51%
Endcaps 0-11% 0-11%
Particle misid. 0-21% 6-37% Y 6-21%
Pair acceptance 15% 15% Y 0%
Event selection 0-7% 0-7% Y 0%
Bin migration negligible negligible negligible
Time dependent effects negligible negligible negligible
Luminosity negligible negligible negligible
Trigger efficiency negligible negligible negligible
Syst. subtotal 15-27% 16-41% 6-21%
Total 15-47% 17-46%
Table 1: Summary of statistical and systematic uncertainties on the correlation functions from various sources, listed separately for deuterium (D) and for the nuclear targets (A). We also list the uncertainties for the ratio $R=C_{A}/C_{D}$, and note whether the systematic uncertainties for the correlation functions are correlated between the nuclear and deuterium targets.
Table 2: Same as Table 4, for the widths, $\sigma$, and broadenings, $b$.
Figure 2: Correlation functions for D, C, Fe, and Pb (from top to bottom, first four rows), with $\Delta\phi$ on the $x$ axis within each panel and different bins of $\Delta Y$ for each column. These are compared to calculations from the BeAGLE model Chang:2022hkt (curves). The horizontal caps in the error bars represent the systematic uncertainties, while the outer bars represent the total systematic and statistical uncertainty (added in quadrature). Bottom row: same for the ratios between the nuclear and deuterium correlation functions.
5 Results and discussion
Figure 2 shows the correlation function $C(\Delta\phi,\Delta Y)$ with different panels representing different $\Delta Y$ bins and the $x$ axis representing $|\Delta\phi|$ [^5] We bin in $|\Delta\phi|$ rather than $\Delta\phi$ since the observed correlation function is symmetrical with respect to $\Delta\phi=\pi$.. We note that the correlation function has a maximum, or “peak [^6] Had we shown $\Delta\phi$ from 0 to $2\pi$, as was done in Ref. PhysRevC.111.035201, this would appear as a symmetrical peak, since the correlation function is symmetrical about $\Delta\phi=\pi$. That is, the correlation function should have the same value at $\Delta\phi=179^{\circ}$ and at $\Delta\phi=181^{\circ}$.”, at $|\Delta\phi|=\pi$ (that is, where the two hadrons are azimuthally opposite one another) in every bin for every target. Further, this peak is tallest within the $1.0<\Delta Y<1.5$ bin (third column). This happens to coincide with the average rapidity separation between the $\gamma^{*}N$ system and the lab frame (about 1.2), which would be expected if the knocked-out proton rapidity distribution were peaked at that of the lab frame, and the leading pion rapidity distribution were peaked at the $\gamma^{*}N$ system. The peak at $|\Delta\phi|=\pi$ is well pronounced for projections where $\Delta Y$ is small, and more spread out in the slices where $\Delta Y$ is large.
The ratios of the correlation functions for nuclei to those of deuterium are shown in the bottom row of Fig. 2 and can reach up to approximately 30 for some bins. We note that these ratios are smaller for azimuthally opposite pairs ($|\Delta\phi|\approx\pi$) than for azimuthally adjacent pairs ($|\Delta\phi|\approx 0$). This is similar to observations in dipion-production measurements CLAS:2022asf, PhysRevC.111.035201, which were attributed to absorption effects coupled with geometrical biases caused by survival selection (that is, hadrons that are absorbed in the nucleus are not detected.) Fialkowski:2007pt.
We further note that, except when $\Delta Y$ is small and $|\Delta\phi|\approx\pi$, the correlation functions are larger for the nuclear targets than for deuterium (as indicated by the ratio, $R$, being greater than one). Since the correlation functions are related to the proton yield per leading pion, $N_{e^{\prime}\pi p}/N_{\pi p}$, (see Eq. 3), it follows that more protons are produced per leading-$\pi^{+}$ event in nuclei than in deuterium. This can partially be explained by the fact that there are more protons that can be knocked out of a larger nucleus than a smaller nucleus. However, this trend appears to saturate, as the values for the correlation functions for iron and lead are within about 30% of one another despite lead having more than three times as many protons as iron. In a scenario where only the protons in or near the path of the cascade could be knocked out, the number of protons that could be knocked out would scale with the nuclear radius, which can be proxied by $\sqrt[3]{A}$, rather than the atomic number, $Z$. In this case, about ($\sqrt[3]{A_{\rm Pb}}/\sqrt[3]{A_{\rm Fe}}=\sqrt[3]{208/56}\approx 1.55$ times as many protons could be knocked out from Pb than from Fe, which is closer to what we observe than if this were scaled with $Z$. Further, the cascade has a limited amount of energy, some of which causes protons to be knocked out and part of which goes into producing other particles, therefore limiting the number of protons that can be knocked out per reaction. This explanation can be tested by comparing these results to future measurements at higher energy, such as a recent experiment with the CLAS12 detector Burkert:2020akg, which used D, C, Al, Cu, Sn, and Pb targets. According to this interpretation, the saturation effect should occur at a larger value of $A$ for higher energies than at lower energies.
We also note that in the highest $\Delta Y$ bin (rightmost column of Fig. 2), the protons and pions are very weakly correlated with one another, as indicated by the correlation function having little-to-no slope with respect to $\Delta\phi$, and that the average number of protons per pion relative to deuterium (see bottom-right panel of Fig. 2) is $\approx$2 for all three targets. Our observation that the protons are very weakly correlated with far-forward pions suggests that future measurements could, by selecting events with large or small numbers of protons, seek to enhance or decrease nuclear effects without biasing the forward pion kinematics RoblesGajardo:2022efe.
We compare our results in Fig. 2 with the calculations of the BeAGLE model Chang:2022hkt (version 1.01.03), which was developed as a benchmark model for the Electron-Ion Collider Accardi:2012qut, AbdulKhalek:2021gbh. The BeAGLE event generator integrates modules from various sources. The primary interaction is managed by PYTHIA6 Sjostrand:2006za. It incorporates nuclear parton distribution functions (PDFs) from LHAPDF5 Whalley:2005nh, intranuclear scattering via DPMJet Roesler:2000he, and the geometric density of nucleons through PyQM Dupre:2011afa. Nuclear remnant de-excitations and decays are processed using FLUKA Ferrari:2005zk. We determined the correlation functions from these simulations by applying Eq. 3 to the number of simulated events within each kinematic bin. The value of $C_{0}$ used for normalizing the simulated correlation functions was determined using the deuterium simulations and is independent of the value used for data.
We find that the trends are similar to those of the data; however, there are numerical discrepancies. Notably, the predicted distribution in $\Delta Y$ appears to be shifted by one bin, or 0.5 rapidity units compared to the data, causing the peak value of the correlation function to be in the 4th panel ($1.5<\Delta Y<2.0$) instead of the 3rd panel ($1.0<\Delta Y<1.5$) as in the data. Further, the predicted curves in a given panel appear to match the data points in the panel to the left of it better than those in the same panel. This could indicate a mis-tuning within the BeAGLE model (possibly in the underlying Pythia6 parameters or in BeAGLE’s modeling of Fermi motion or other phenomena), causing either the generated final-state pions to tend towards higher rapidity than those in the data, or the generated protons to tend towards lower rapidity than those in the data.
Figure 3: Top row: RMS width $\sigma$ as a function of $\sqrt[3]{A}$ (left-most panel), and as a function of $\Delta Y$ for each target (other panels). Bottom row: same for the azimuthal broadening, $b$. The horizontal caps in the error bars represent the systematic uncertainties, while the outer bars represent the total systematic and statistical uncertainty (added in quadrature). The results are compared to the predictions from the BeAGLE Chang:2022hkt, eHIJING PhysRevD.110.034001, and GiBUU Buss:2011mx models, as well as analogous measurements of the dipion channel from Ref. PhysRevC.111.035201 (gray, open squares). The broadenings for the dipion channel were reported in Ref. PhysRevC.111.035201 with respect to deuterium, and have been re-calculated with respect to carbon in order to compare them to the results of this work.
From the correlation functions, we determined the RMS widths, $\sigma$, integrated in $\Delta Y$, which we show in the top left panel of Fig. 3 as a function of $\sqrt[3]{A}$ (which proxies the radius of the nucleus). We note that $\sigma$ increases monotonically with respect to increasing nuclear size, with values of about 1.13, 1.32, 1.41, and 1.45 for D, C, Fe, and Pb, respectively. This increase is further reflected in the azimuthal broadenings, $b$, which increase as a function of $A$, as shown in the lower left panel of the same figure. This trend is similar to what was observed in the dipion measurements PhysRevC.111.035201 (gray open squares in Fig. 3). We note that this increase in width with respect to nuclear size is steeper in this work than in Ref. PhysRevC.111.035201. This seems to suggest that low-energy protons undergo stronger final-state interactions than secondary pions. In the context of some models, such as the GiBUU model Buss:2011mx, when secondary pions are produced, they have a “pre-hadronic” phase when they are color neutral but are not fully formed. During this phase, the “pre-hadrons” have a weaker interaction with the nuclear material than after they have fully formed. Since protons from knockout are already fully formed, they do not have this pre-hadronic phase; therefore they interact with the nucleus as soon as they are struck by the cascade. Another possible explanation is that the protons have lower kinetic energy than pions of the same momentum, and are therefore more susceptible to energy loss and absorption effects.
In the subsequent columns of this figure, we show the widths and broadenings as a function of $\Delta Y$ for each target. The widths increase monotonically as a function of $\Delta Y$ (which was also observed for the dipion channel in Ref. PhysRevC.111.035201, however we observe that the slope of this dependence is larger for the $\pi p$ channel than for the dipion channel). The broadening is consistently negative (up to $\approx$-0.8) for deuterium and positive for iron (up to $\approx 0.6$) and lead (up to $\approx 0.7$), with the exception of one bin in which the uncertainties are very large.
These trends in the widths and broadenings are reproduced by the BeAGLE model, which predicts the values of $\sigma$ within about one error bar for the integrated results (upper left panel of Fig. 3). However, the agreement with the data is not as good within slices in $\Delta Y$ as in the integrated results.
We also compared our results for the widths and broadenings to the predictions of the eHIJING model PhysRevD.110.034001 (version 1.0) and the GiBUU model Buss:2011mx (2021 release) in Fig. 3. We also provide a comparison between these models’ predictions for the correlation functions and the ratio, $R$, in the Supplementary Materials supplementary. In eHIJING, the hard-scattering process, along with initial- and final-state fragmentation, is simulated using PYTHIA8 bierlich2022comprehensive. The model incorporates nuclear modifications to PDFs, fragmentation, and parton showering. It also handles multiple collisions between partons and the rest of the nucleus, with a cross section proportional to the transverse-momentum-dependent gluon distribution density. Like the BeAGLE model, eHIJING was developed for the Electron-Ion Collider.
The GiBUU model incorporates treatment of final-state interactions, absorption, and production mechanisms with elastic and inelastic channels. The GiBUU model describes the single-hadron and di-hadron data reasonably well from CLAS CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201 and HERMES HERMES:2003icw, Airapetian:2007vu, Airapetian:2011jp. We modified the default values of several parameters related to parton distributions and fragmentation in free nucleons, based on a tuning study that used an independent proton-target dataset from Ref. CLAS:2022sqt, and which was further validated with dipion correlations in Ref. PhysRevC.111.035201. This tuning does not modify the nuclear effects.
Both the eHIJING and GiBUU models reproduce the trends of the data. The eHIJING predictions have a smaller variation of $\sigma$ with respect to both nuclear size and $\Delta Y$. Further, it underestimates the magnitude of the broadening compared to the data. The GiBUU model predicts the widths very well for deuterium (both in the $\Delta Y$ bins and when integrated in $\Delta Y$), but underpredicts the widths for the nuclear targets.
6 Summary and conclusions
We have reported the first measurement of angular correlations between leading pions and protons emerging from nuclear fragmentation, opening a new avenue to study how nuclei respond to fast hadrons.
Using electron–nucleus scattering data collected with the CLAS detector on deuterium and a broad set of heavier targets, we measured correlation functions that peak when the two particles are azimuthally back‑to‑back and their rapidity difference matches that between the virtual-photon+struck-nucleon frame and the initial‑nucleus rest frame. As the rapidity separation increases, the azimuthal peak broadens; this broadening is more pronounced in larger nuclei.
The azimuthal‑peak width grows systematically with nuclear size, and the proton yield per pion also rises but it appears to saturate for the heaviest targets. These patterns are consistent with a picture in which hadron absorption and interactions in the hadronic cascade widens the correlation and the larger pool of protons in heavier nuclei enhances the yield.
Compared with earlier dipion measurements PhysRevC.111.035201, the $\pi p$ channel shows a markedly stronger dependence on both nuclear size and rapidity separation. The trends are qualitatively reproduced by the state-of-the-art $eA$ event-generator models BeAGLE, eHIJING, and GiBUU, but quantitative differences remain, highlighting opportunities to refine the treatment of nuclear fragmentation.
Our results establish a benchmark for forthcoming di-hadron and slow-nucleon production studies in DIS, as well as for the development of corresponding event-generator models for experiments at Jefferson Lab Burkert:2018nvj, Burkert:2020akg, Arrington:2021alx and at future Electron–Ion Colliders Accardi:2012qut, Anderle:2021wcy.
7 Code and Data Availability
We provide in Ref. zenodoRivet the code we used to extract the $(\Delta\phi,\Delta Y)$ histograms from the output files of event generators (in order to produce the correlation functions) in the form of a module in the Rivet framework 10.21468/SciPostPhys.8.2.026.
The values of the correlation functions, the ratio $R$, RMS widths, and broadenings reported in this work can be found in the Supplementary Materials supplementary.
8 Acknowledgements
The authors acknowledge the staff of the Accelerator and Physics Divisions at the Thomas Jefferson National Accelerator Facility who made this experiment possible. We thank Kai Gallmeister for help in setting up the GiBUU event generator and Weiyao Ke for help in setting up the eHIJING event generator. This work is supported by the Chilean Agencia Nacional de Investigacion y Desarollo (ANID), FONDECYT Grants No. 1221827 and No. 1240904, ANID PIA/APOYO AFB230003 and by the ANID-Millennium Science Initiative Program–ICN2019_044, by the U.S. Department of Energy, the Italian Instituto Nazionale di Fisica Nucleare, the French Centre National de la Recherche Scientifique, the French Commissariat à l’Energie Atomique, the United Kingdom Science and Technology Facilities Council (STFC), the Scottish Universities Physics Alliance (SUPA), the National Research Foundation of Korea, the National Science Foundation (NSF), the HelmholtzForschungsakademie Hessen für FAIR (HFHF), the Ministry of Science and Higher Education of the Russian Federation, and the Office of Research and Economic Development at Mississippi State University. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 804480). The Southeastern Universities Research Association operates the Thomas Jefferson National Accelerator Facility for the United States Department of Energy under Contract No. DE-AC05-06OR23177.

### v2

@@
 Abstract
-We report on the first measurement of angular correlations between high‑energy pions and slow protons in electron–nucleus ($eA$) scattering, providing a new probe of how a nucleus responds to a fast‑moving quark.
+We report on the first measurement of angular correlations between high‑energy pions and slow protons in electron–nucleus (eAeA) scattering, providing a new probe of how a nucleus responds to a fast‑moving quark.
 The experiment employed the CLAS detector with a 5-GeV electron beam incident on deuterium, carbon, iron, and lead targets.
@@
 The proton‑to‑pion yield ratio likewise rises with nuclear mass, although the increase appears to saturate for the heaviest targets.
-These trends are qualitatively reproduced by state‑of‑the‑art $eA$ event generators, including BeAGLE, eHIJING, and GiBUU, indicating that current descriptions of target fragmentation rest on sound theoretical footing.
-At the same time, the precision of our data exposes model–dependent discrepancies, delineating a clear path for future improvements in the treatment of cold-nuclear matter effects in $eA$ scattering.
+These trends are qualitatively reproduced by state‑of‑the‑art eAeA event generators, including BeAGLE, eHIJING, and GiBUU, indicating that current descriptions of target fragmentation rest on sound theoretical footing.
+At the same time, the precision of our data exposes model–dependent discrepancies, delineating a clear path for future improvements in the treatment of cold-nuclear matter effects in eAeA scattering.
 1 Introduction
@@
 The present work addresses this gap by employing pion-proton correlations over a wide azimuthal and rapidity range, thereby linking studies of current and nuclear fragmentation in DIS off deuterium and heavy nuclei.
-In this study, we examine electron–nucleus ($eA$) DIS events that contain a leading $\pi^{+}$ accompanied by a secondary proton, as illustrated in Fig.
+In this study, we examine electron–nucleus (eAeA) DIS events that contain a leading $\pi^{+}$ accompanied by a secondary proton, as illustrated in Fig.
 1.
@@
 Larionov:2018igy showed that slow-neutrons provide similar insight into intranuclear cascades.
-The study of slow protons (or more generally, slow nucleons) will also play an important role in the future Electron-Ion Collider Accardi:2012qut, AbdulKhalek:2021gbh, since slow nucleons will be measured with dedicated detectors, to enable various studies such as spectator tagging in electron-deuteron collisions, centrality estimations, and hadronization studies in $eA$ collisions AbdulKhalek:2021gbh, Chang:2022hkt, RoblesGajardo:2022efe.
+The study of slow protons (or more generally, slow nucleons) will also play an important role in the future Electron-Ion Collider Accardi:2012qut, AbdulKhalek:2021gbh, since slow nucleons will be measured with dedicated detectors, to enable various studies such as spectator tagging in electron-deuteron collisions, centrality estimations, and hadronization studies in eAeA collisions AbdulKhalek:2021gbh, Chang:2022hkt, RoblesGajardo:2022efe.
 2 Experimental setup
 […]
-The CLAS CLAS:2003umf detector was based on a six-fold symmetric toroidal magnet, which defined six sectors instrumented with drift chambers (DC), time-of-flight scintillation counters (TOF), Cherenkov counters (CC), and an electromagnetic calorimeter (EC).
+The CLAS CLAS:2003umf detector was based on a six-fold symmetric toroidal magnet, which defined six sectors instrumented with drift chambers (DC), time-of-flight scintillation counters (TOF), Cherenkov counters ($C$), and an electromagnetic calorimeter (EC).
 Its polar angular acceptance ranged from 8 to 140 for the drift chambers and TOF, and from 8 to 45 for the Cherenkov counters and calorimeter.
@@
 CLAS:2012tlh, CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201, electrons were identified by matching negatively charged tracks measured in the DC with hits in the TOF and EC.
-Background from $\pi^{-}$ was suppressed to the $1\%$ level using the CC and the EC.
+Background from $\pi^{-}$ was suppressed to the $1\%$ level using the $C$ and the EC.
 Charged pions and protons were identified with TOF hits consistent with those types of hadrons with momentum determined in the DCs.
-The selection for $\pi^{+}$ with momentum above 2.7 GeV/c/c was further refined with the CC to suppress proton background.
+The selection for $\pi^{+}$ with momentum above 2.7 GeV/c/c was further refined with the $C$ to suppress proton background.
 Fiducial cuts on momentum and angles were used in order to avoid regions with steeply varying acceptance or low resolution.
@@
 This explanation can be tested by comparing these results to future measurements at higher energy, such as a recent experiment with the CLAS12 detector Burkert:2020akg, which used D, C, Al, Cu, Sn, and Pb targets.
-According to this interpretation, the saturation effect should occur at a larger value of $A$ for higher energies than at lower energies.
+According to this interpretation, the saturation effect should occur at a larger value of AA for higher energies than at lower energies.
 We also note that in the highest $\Delta Y$ bin (rightmost column of Fig.
@@
 We note that $\sigma$ increases monotonically with respect to increasing nuclear size, with values of about 1.13, 1.32, 1.41, and 1.45 for D, C, Fe, and Pb, respectively.
-This increase is further reflected in the azimuthal broadenings, $b$, which increase as a function of $A$, as shown in the lower left panel of the same figure.
+This increase is further reflected in the azimuthal broadenings, $b$, which increase as a function of AA, as shown in the lower left panel of the same figure.
 This trend is similar to what was observed in the dipion measurements PhysRevC.111.035201 (gray open squares in Fig.
@@
 Compared with earlier dipion measurements PhysRevC.111.035201, the $\pi p$ channel shows a markedly stronger dependence on both nuclear size and rapidity separation.
-The trends are qualitatively reproduced by the state-of-the-art $eA$ event-generator models BeAGLE, eHIJING, and GiBUU, but quantitative differences remain, highlighting opportunities to refine the treatment of nuclear fragmentation.
+The trends are qualitatively reproduced by the state-of-the-art eAeA event-generator models BeAGLE, eHIJING, and GiBUU, but quantitative differences remain, highlighting opportunities to refine the treatment of nuclear fragmentation.
 Our results establish a benchmark for forthcoming di-hadron and slow-nucleon production studies in DIS, as well as for the development of corresponding event-generator models for experiments at Jefferson Lab Burkert:2018nvj, Burkert:2020akg, Arrington:2021alx and at future Electron–Ion Colliders Accardi:2012qut, Anderle:2021wcy.

### v3

@@
 Abstract
-We report on the first measurement of angular correlations between high‑energy pions and slow protons in electron–nucleus (eAeA) scattering, providing a new probe of how a nucleus responds to a fast‑moving quark.
+We report on the first measurement of angular correlations between high‑energy pions and slow protons in electron–nucleus ($eA$) scattering, providing a new probe of how a nucleus responds to a fast‑moving quark.
 The experiment employed the CLAS detector with a 5-GeV electron beam incident on deuterium, carbon, iron, and lead targets.
@@
 The proton‑to‑pion yield ratio likewise rises with nuclear mass, although the increase appears to saturate for the heaviest targets.
-These trends are qualitatively reproduced by state‑of‑the‑art eAeA event generators, including BeAGLE, eHIJING, and GiBUU, indicating that current descriptions of target fragmentation rest on sound theoretical footing.
-At the same time, the precision of our data exposes model–dependent discrepancies, delineating a clear path for future improvements in the treatment of cold-nuclear matter effects in eAeA scattering.
+These trends are qualitatively reproduced by state‑of‑the‑art $eA$ event generators, including BeAGLE, eHIJING, and GiBUU, indicating that current descriptions of target fragmentation rest on sound theoretical footing.
+At the same time, the precision of our data exposes model–dependent discrepancies, delineating a clear path for future improvements in the treatment of cold-nuclear matter effects in $eA$ scattering.
 1 Introduction
 The production and propagation of hadrons through nuclei remain central topics in quantum chromodynamics (QCD).
-Key questions include: “How does the process of hadronization manifest within a nucleus?”, and “How are different hadrons produced in a single scattering event correlated with each other?” LRP2023.
+Key questions include: “How does the process of hadronization manifest within a nucleus?”, and “How are different hadrons produced in a single scattering event correlated with each other?” Nuclear Science Advisory Committee .
 Addressing these questions calls for detailed studies of diverse reaction channels, encompassing multiple hadron flavor combinations in the final state and spanning a broad kinematic range, to probe the different timescales of hadronization.
-Despite extensive theoretical and phenomenological work on nucleon production in nuclear deep-inelastic scattering (DIS) E665:1994aiu, Degtyarenko:1997, BEBCWA59:1989ayi, CiofidegliAtti:2004pv, Palli:2009it, Larionov:2018igy, PhysRevD.100.073010, Chang:2022hkt, RoblesGajardo:2022efe, no dedicated study has correlated the production of low-energy protons with the kinematics of leading hadrons, whose spectra are known to be modified by cold nuclear matter effects as observed by the HERMES HERMES:2000ytc, HERMES:2003icw, Airapetian:2007vu, Airapetian:2009jy, Airapetian:2011jp and CLAS CLAS:2021jhm, CLAS:2022oux, CLAS:2012tlh, kbhz-h4jv collaborations.
+Despite extensive theoretical and phenomenological work on nucleon production in nuclear deep-inelastic scattering (DIS) Adams and others [1995], Degtyarenko et al.
+[1997], Matsinos and others [1989], Ciofi degli Atti and Kopeliovich [2005], Palli et al.
+[2009], Larionov and Strikman [2020], Cai et al.
+[2019], Chang et al.
+[2022], Robles Gajardo et al.
+[2022], no dedicated study has correlated the production of low-energy protons with the kinematics of leading hadrons, whose spectra are known to be modified by cold nuclear matter effects as observed by the HERMES Airapetian and others [2001, 2003, 2007, 2010, 2011] and CLAS Moran and others [2022], Chetry and others [2023], El Fassi and others [2012], Mineeva et al.
+[2025] collaborations.
 The present work addresses this gap by employing pion-proton correlations over a wide azimuthal and rapidity range, thereby linking studies of current and nuclear fragmentation in DIS off deuterium and heavy nuclei.
-In this study, we examine electron–nucleus (eAeA) DIS events that contain a leading $\pi^{+}$ accompanied by a secondary proton, as illustrated in Fig.
+In this study, we examine electron–nucleus ($eA$) DIS events that contain a leading $\pi^{+}$ (which retains most of the energy transferred to the struck quark) accompanied by a secondary proton, as illustrated in Fig.
 1.
@@
 Most of the observed protons come from fragmentation of the nuclear remnant.
-Because the proton’s rest energy is already present in the initial state, a smaller amount of energy from the hadronic cascade is necessary to knock a proton out of a nucleus than produce a secondary pion with the same momentum.
-This allows the current measurement to probe lower energy scales of the hadronization process and transport properties of the nucleus than in existing dipion measurements PhysRevC.111.035201.
+Because the proton’s rest energy is already present in the initial state, a smaller amount of energy from the fast quark or its hadronic cascade is necessary to knock a proton out of a nucleus than produce a secondary pion with the same momentum.
+This allows the current measurement to probe lower energy scales of the hadronization process and transport properties of the nucleus than in existing dipion measurements Paul and others [2025].
 Figure 1: Illustration of nuclear deep‑inelastic scattering featuring a leading charged pion and secondary protons in the final state.
-Protons produced through nuclear fragmentation, historically referred to as “grey” protons, were investigated in earlier scattering experiments with both hadronic and leptonic beams E910:1999ozb, E665:1994aiu, Degtyarenko:1997.
+Protons produced through nuclear fragmentation, historically referred to as “grey” protons, were investigated in earlier scattering experiments with both hadronic and leptonic beams Chemakin and others [1999], Adams and others [1995], Degtyarenko et al.
+[1997].
 Measurements of slow protons, together with model calculations, served to estimate the collision centrality, expressed as the mean number of projectile–nucleon interactions.
 Lepton measurements offer a cleaner test of intra-nuclear hadron-cascade models; in hadron collisions, multiple projectile–nucleon interactions complicate particle‐production mechanisms.
-Similar studies were carried out with neutrino DIS on nuclear targets BEBCWA59:1989ayi, although with much more limited precision.
+Similar studies were carried out with neutrino DIS on nuclear targets Matsinos and others [1989], although with much more limited precision.
 Ciofi degli Atti et al.
-CiofidegliAtti:2004pv, Palli:2009it proposed that the dependence of slow-proton production over a wide kinematic range can serve to discriminate among hadronization models, and their calculation reproduced well the data from Ref.
-E665:1994aiu.
+Ciofi degli Atti and Kopeliovich [2005], Palli et al.
+[2009] proposed that the dependence of slow-proton production over a wide kinematic range can serve to discriminate among hadronization models, and their calculation reproduced well the data from Ref.
+Adams and others [1995].
 Larionov et al.
-Larionov:2018igy showed that slow-neutrons provide similar insight into intranuclear cascades.
-The study of slow protons (or more generally, slow nucleons) will also play an important role in the future Electron-Ion Collider Accardi:2012qut, AbdulKhalek:2021gbh, since slow nucleons will be measured with dedicated detectors, to enable various studies such as spectator tagging in electron-deuteron collisions, centrality estimations, and hadronization studies in eAeA collisions AbdulKhalek:2021gbh, Chang:2022hkt, RoblesGajardo:2022efe.
+Larionov and Strikman [2020] showed that slow-neutrons provide similar insight into intranuclear cascades.
+The study of slow protons (or more generally, slow nucleons) will also play an important role in the future Electron-Ion Collider Accardi and others [2016], Abdul Khalek and others [2022], since slow nucleons will be measured with dedicated detectors, to enable various studies such as spectator tagging in electron-deuteron collisions, centrality estimations, and hadronization studies in $eA$ collisions Abdul Khalek and others [2022], Chang et al.
+[2022], Robles Gajardo et al.
+[2022].
+Further, the measurement of $\pi p$ correlations in $eA$ DIS can be used as a reference for analogous neutrino measurements, for instance at MINERvA Aliaga and others [2014], which overlaps the kinematics of the measurements presented in this work, or with legacy data from NOvA Ayres et al.
+[2007].
+It was suggested in Ref.
+Cai et al.
+[2019] that left-right asymmetries in $\pi p$ correlations could be measured in $\nu A$ scattering to probe pion-absorption effects.
 2 Experimental setup
 […]
-The CLAS CLAS:2003umf detector was based on a six-fold symmetric toroidal magnet, which defined six sectors instrumented with drift chambers (DC), time-of-flight scintillation counters (TOF), Cherenkov counters ($C$), and an electromagnetic calorimeter (EC).
+The CLAS Mecking and others [2003] detector was based on a six-fold symmetric toroidal magnet, which defined six sectors instrumented with drift chambers (DC), time-of-flight scintillation counters (TOF), Cherenkov counters (CC), and an electromagnetic calorimeter (EC).
 Its polar angular acceptance ranged from 8 to 140 for the drift chambers and TOF, and from 8 to 45 for the Cherenkov counters and calorimeter.
@@
 Following Refs.
-CLAS:2012tlh, CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201, electrons were identified by matching negatively charged tracks measured in the DC with hits in the TOF and EC.
-Background from $\pi^{-}$ was suppressed to the $1\%$ level using the $C$ and the EC.
+El Fassi and others [2012], Moran and others [2022], Paul and others [2022, 2025], electrons were identified by matching negatively charged tracks measured in the DC with hits in the TOF and EC.
+Background from $\pi^{-}$ was suppressed to the $1\%$ level using the CC and the EC.
 Charged pions and protons were identified with TOF hits consistent with those types of hadrons with momentum determined in the DCs.
-The selection for $\pi^{+}$ with momentum above 2.7 GeV/c/c was further refined with the $C$ to suppress proton background.
+The selection for $\pi^{+}$ with momentum above 2.7 GeV/c/c was further refined with the CC to suppress proton background.
 Fiducial cuts on momentum and angles were used in order to avoid regions with steeply varying acceptance or low resolution.
 The protons’ energies were corrected for energy loss in the target using the same procedure as was used in Ref.
-CLAS:2022oux.
+Chetry and others [2023].
 Similar to Refs.
-CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201, we selected events with $Q^{2}>1$ $2W>2 GeV/c2c^{2}$ and $2.3<\nu<4.2$ GeV.
+Moran and others [2022], Paul and others [2022, 2025], we selected events with $Q^{2}>1$ $2W>2 GeV/c2c^{2}$ and $2.3<\nu<4.2$ GeV.
 Here, $Q^{2}$ is the square of the four-momentum transfer from the electron to the struck nucleon,
@@
 is the magnitude of the four-momentum of the virtual photon+struck nucleon system, where $m_{p}$ is the mass of a proton, $\nu=E-E^{\prime}$ is the energy transfer, and $E$ and $E^{\prime}$ are the beam- and scattered-electron energies.
-We selected events with a “leading” pion, defined as having fractional energy $z_{1}=E_{h}/\nu>0.5$ HERMES:2005mar, CLAS:2022asf, PhysRevC.111.035201, where $E_{h}$ is the energy of the pion.
-Following Refs.
-CLAS:2022oux, Hen:2014nza, we required the proton to have momentum greater than 350 MeV/c/c so that it could be well-reconstructed.
+The cuts on $Q^{2}$ and $W$ are conventional for DIS measurements.
+The minimum cut on $\nu$ is just above the minimum value of $\nu$ possible with these cuts on $Q^{2}$ and $W$.
+The maximum value of $\nu$ corresponds to a cut on the inelasticity, $y\equiv\nu/E<0.84$.
+We selected events with a “leading” pion, defined as having fractional energy $z_{1}=E_{h}/\nu>0.5$ Airapetian and others [2006], Paul and others [2022, 2025], where $E_{h}$ is the energy of the pion.
+The purpose of this cut is to increase the likelihood that the pion originated from the struck quark and to avoid ambiguity on which pion in the event is considered “leading”.
+Following Ref.
+Hen and others [2014], we required the proton to have momentum greater than 350 MeV/c/c, since the proton detection efficiency drops off below this threshold.
 The maximum proton momentum in our accepted $\pi p$ event sample was about 2.7 GeV/c/c.
-In order to ensure that the azimuthal angles of both particles are well defined, we further required that both the proton and the leading pion have transverse momenta (defined with respect to the momentum-transfer vector $\vec{q}$) greater than 70 MeV/c/c.
+In order to ensure that the azimuthal angles of both particles can be reconstructed with reasonable resolution, we further required that both the proton and the leading pion have transverse momenta (defined with respect to the momentum-transfer vector $\vec{q}$) greater than 70 MeV/c/c.
 We selected events within the range $0.0<\Delta Y<3.0$, where $\Delta Y\equiv Y_{\pi}-Y_{p}$ is the rapidity separation between the pion and the proton.
@@
 where $E_{h}$ is the hadron’s energy and $p_{z,h}$ is its momentum in the momentum-transfer direction.
+This range in $\Delta Y$ contains nearly 99% of all $\pi^{+}p$ pairs, and allows for a convenient for binning in $\Delta Y$.
+This differs from the $\Delta Y$ range in the di-pion analysis in Ref.
+Paul and others [2025], where the observed range in $\Delta Y$ for $\pi^{+}\pi^{-}$ was approximately -0.5 to 2.5.
 We selected particles arising from scattering from either the deuterium or nuclear targets by using the longitudinal vertex position defined by intersecting their trajectories with the beamline.
-The resulting vertex resolution ensured negligible ambiguity in the target tagging CLAS:2021jhm.
+The resulting vertex resolution ensured negligible ambiguity in the target tagging Moran and others [2022].
 We used the electron, the leading pion, and the proton variables to measure the 2-dimensional correlation function, defined as
@@
 Following Ref.
-PhysRevC.111.035201, we also used a data-driven method to correct for contamination of the deuterium sample due to scattering off the aluminum endcap walls of the liquid deuterium cells.
-Details for both of these corrections can be found in the Supplementary Materials supplementary.
+Paul and others [2025], we also used a data-driven method to correct for contamination of the deuterium sample due to scattering off the aluminum endcap walls of the liquid deuterium cells.
+Details for both of these corrections can be found in the Supplementary Materials 1.
 The normalization factor $C_{0}$ was chosen such that the integral of $C(\Delta\phi,\Delta Y)$ over $\Delta\phi$ and $\Delta Y$ equals unity for deuterium.
@@
 Explicitly, the formula we used was
-$C_{0}=\frac{N^{D}_{e^{\prime}\pi}}{N^{D,\rm tot}_{e^{\prime}\pi p}}$ (4)
+$C_{0}=\frac{N^{D}_{e^{\prime}\pi}}{N^{D,\rm tot}_{e^{\prime}\pi p}},$ (4)
 where $N^{D}_{e^{\prime}\pi}$ is the number of events from the deuterium dataset with the scattered electron and leading pion, and $N^{D,\rm tot}_{e^{\prime}\pi p}$ is the total number of events in the deuterium dataset with the scattered electron, leading pion, and proton, for all kinematic bins.
+The purpose of this normalization is to make the correlation function an observable that is insensitive to the absolute scale of the single-proton efficiency of the detector systems.
 From here, we determined the nuclear-to-deuterium ratio of the correlation functions, $R=C_{A}/C_{D}$, which can be used to compare the correlation functions of the nuclear targets with those of deuterium.
-We determined from these correlation functions the azimuthal width $\sigma$, defined by PhysRevC.111.035201
+Since the same $C_{0}$ value is used for nuclear and deuterium targets, it cancels out in the ratio, which is then equivalent to
+$R\equiv\frac{N^{A}_{e^{\prime}\pi^{+}p}/N^{A}_{e^{\prime}\pi^{+}}}{N^{D}_{e^{\prime}\pi^{+}p}/N^{D}_{e^{\prime}\pi^{+}}}.$ (5)
+We determined from these correlation functions the azimuthal width $\sigma$, defined by Paul and others [2025]
 $\sigma=\sqrt{\frac{\sum\limits_{i\in\rm bins}C_{i}\left(\Delta\phi_{i}-\pi\right)^{2}}{\sum\limits_{i\in\rm bins}C_{i}}},$ (6)
-as well as the azimuthal broadening PhysRevC.111.035201 (using carbon as our reference target [^2] We chose carbon instead of deuterium (as was done in Ref.
-PhysRevC.111.035201) for the reference target here, since the results for the nuclear targets are more similar to one another than they are to deuterium.),
+as well as the azimuthal broadening Paul and others [2025] (using carbon as our reference target [^2] We chose carbon instead of deuterium (as was done in Ref.
+Paul and others [2025]) for the reference target here, since the results for the nuclear targets are more similar to one another than they are to deuterium.),
 $b=\pm\sqrt{\left|\sigma_{A}^{2}-\sigma_{C}^{2}\right|},$ (7)
@@
 We performed studies on various possible sources of systematic uncertainties using data and simulation studies.
-For the simulation studies we used the PYTHIA6 event generator and the GSIM package GSIM, which is based on GEANT3 Brun:1994aa [^3] It should be noted that Geant4 GEANT4:2002zbu had not been released yet when GSIM was being developed, and that GEANT3 was still the standard at the time that the data were taken., to simulate the response of the CLAS detector and dual-target setup Hakobyan:2008kua.
+For the simulation studies we used the PYTHIA6 event generator and the GSIM package Wolin [1996], which is based on GEANT3 Brun et al.
+[1994] [^3] It should be noted that Geant4 Agostinelli and others [2003] had not been released yet when GSIM was being developed, and that GEANT3 was still the standard at the time that the data were taken., to simulate the response of the CLAS detector and dual-target setup Hakobyan [2008].
 The simulation was tuned to provide a reasonable description of the data.
@@
 The reason for this is that the measured $\sigma$ values for C and Fe are very close together, much less than one statistical error bar.
-Therefore, the uncertainties on their quadrature difference are very large..
-Details of these studies on systematic uncertainties can be found in the Supplementary Materials supplementary.
+Therefore, the relative uncertainties on their quadrature difference are very large..
+Details of these studies on systematic uncertainties can be found in the Supplementary Materials 1.
 Source $\Delta C/C$ (D) $\Delta C/C$ (A) corr. D vs A $\Delta R/R$
@@
 Figure 2: Correlation functions for D, C, Fe, and Pb (from top to bottom, first four rows), with $\Delta\phi$ on the $x$ axis within each panel and different bins of $\Delta Y$ for each column.
-These are compared to calculations from the BeAGLE model Chang:2022hkt (curves).
+These are compared to calculations from the BeAGLE model Chang et al.
+[2022] (curves).
 The horizontal caps in the error bars represent the systematic uncertainties, while the outer bars represent the total systematic and statistical uncertainty (added in quadrature).
@@
 We note that the correlation function has a maximum, or “peak [^6] Had we shown $\Delta\phi$ from 0 to $2\pi$, as was done in Ref.
-PhysRevC.111.035201, this would appear as a symmetrical peak, since the correlation function is symmetrical about $\Delta\phi=\pi$.
+Paul and others [2025], this would appear as a symmetrical peak, since the correlation function is symmetrical about $\Delta\phi=\pi$.
 That is, the correlation function should have the same value at $\Delta\phi=179^{\circ}$ and at $\Delta\phi=181^{\circ}$.”, at $|\Delta\phi|=\pi$ (that is, where the two hadrons are azimuthally opposite one another) in every bin for every target.
 Further, this peak is tallest within the $1.0<\Delta Y<1.5$ bin (third column).
-This happens to coincide with the average rapidity separation between the $\gamma^{*}N$ system and the lab frame (about 1.2), which would be expected if the knocked-out proton rapidity distribution were peaked at that of the lab frame, and the leading pion rapidity distribution were peaked at the $\gamma^{*}N$ system.
+This happens to coincide with the average rapidity separation between the $\gamma^{*}N$ system and the lab frame (about 1.2), which would be expected if the knocked-out proton rapidity distribution were peaked at that of the lab frame (as would be the case if their initial momentum distribution is isotropic), and the leading pion rapidity distribution were peaked at the $\gamma^{*}N$ system from which it originated.
+This interpretation could be tested by seeing if this observation continues to hold true in future measurements, e.g in experiments with the CLAS12 detector Burkert and others [2020], where a larger range in $\nu$ and $Q^{2}$ allows a larger range of values for the rapidity of the $\gamma^{*}N$ system.
 The peak at $|\Delta\phi|=\pi$ is well pronounced for projections where $\Delta Y$ is small, and more spread out in the slices where $\Delta Y$ is large.
 The ratios of the correlation functions for nuclei to those of deuterium are shown in the bottom row of Fig.
-2 and can reach up to approximately 30 for some bins.
+2.
+The values of these ratios range from near unity to about 30, depending on the target and kinematic bin.
 We note that these ratios are smaller for azimuthally opposite pairs ($|\Delta\phi|\approx\pi$) than for azimuthally adjacent pairs ($|\Delta\phi|\approx 0$).
-This is similar to observations in dipion-production measurements CLAS:2022asf, PhysRevC.111.035201, which were attributed to absorption effects coupled with geometrical biases caused by survival selection (that is, hadrons that are absorbed in the nucleus are not detected.) Fialkowski:2007pt.
+This is similar to observations in dipion-production measurements Paul and others [2022, 2025], which were attributed to absorption effects coupled with geometrical biases caused by survival selection (that is, hadrons that are absorbed in the nucleus are not detected.) Fialkowski and Wit [2007].
 We further note that, except when $\Delta Y$ is small and $|\Delta\phi|\approx\pi$, the correlation functions are larger for the nuclear targets than for deuterium (as indicated by the ratio, $R$, being greater than one).
-Since the correlation functions are related to the proton yield per leading pion, $N_{e^{\prime}\pi p}/N_{\pi p}$, (see Eq.
+Since the correlation functions are related to the proton yield per leading pion, $N_{e^{\prime}\pi p}/N_{e^{\prime}\pi}$, (see Eq.
 3), it follows that more protons are produced per leading-$\pi^{+}$ event in nuclei than in deuterium.
@@
 In a scenario where only the protons in or near the path of the cascade could be knocked out, the number of protons that could be knocked out would scale with the nuclear radius, which can be proxied by $\sqrt[3]{A}$, rather than the atomic number, $Z$.
-In this case, about ($\sqrt[3]{A_{\rm Pb}}/\sqrt[3]{A_{\rm Fe}}=\sqrt[3]{208/56}\approx 1.55$ times as many protons could be knocked out from Pb than from Fe, which is closer to what we observe than if this were scaled with $Z$.
+In this case, about $\sqrt[3]{A_{\rm Pb}}/\sqrt[3]{A_{\rm Fe}}=\sqrt[3]{208/56}\approx 1.55$ times as many protons could be knocked out from Pb than from Fe, which is closer to what we observe than if this were scaled with $Z$.
 Further, the cascade has a limited amount of energy, some of which causes protons to be knocked out and part of which goes into producing other particles, therefore limiting the number of protons that can be knocked out per reaction.
-This explanation can be tested by comparing these results to future measurements at higher energy, such as a recent experiment with the CLAS12 detector Burkert:2020akg, which used D, C, Al, Cu, Sn, and Pb targets.
-According to this interpretation, the saturation effect should occur at a larger value of AA for higher energies than at lower energies.
+This explanation can be tested by comparing these results to future measurements at higher energy, such as a recent experiment with the CLAS12 detector Burkert and others [2020], which used D, C, Al, Cu, Sn, and Pb targets.
+According to this interpretation, the saturation effect should occur at a larger value of $A$ for higher energies than at lower energies.
 We also note that in the highest $\Delta Y$ bin (rightmost column of Fig.
@@
 2) is $\approx$2 for all three targets.
-Our observation that the protons are very weakly correlated with far-forward pions suggests that future measurements could, by selecting events with large or small numbers of protons, seek to enhance or decrease nuclear effects without biasing the forward pion kinematics RoblesGajardo:2022efe.
+Our observation that the protons are very weakly correlated with far-forward pions suggests that future measurements could, by selecting events with large or small numbers of protons, seek to enhance or decrease nuclear effects without biasing the forward pion kinematics Robles Gajardo et al.
+[2022].
 We compare our results in Fig.
-2 with the calculations of the BeAGLE model Chang:2022hkt (version 1.01.03), which was developed as a benchmark model for the Electron-Ion Collider Accardi:2012qut, AbdulKhalek:2021gbh.
+2 with the calculations of the BeAGLE model Chang et al.
+[2022] (version 1.01.03), which was developed as a benchmark model for the Electron-Ion Collider Accardi and others [2016], Abdul Khalek and others [2022].
 The BeAGLE event generator integrates modules from various sources.
-The primary interaction is managed by PYTHIA6 Sjostrand:2006za.
-It incorporates nuclear parton distribution functions (PDFs) from LHAPDF5 Whalley:2005nh, intranuclear scattering via DPMJet Roesler:2000he, and the geometric density of nucleons through PyQM Dupre:2011afa.
-Nuclear remnant de-excitations and decays are processed using FLUKA Ferrari:2005zk.
+The primary interaction is managed by PYTHIA6 Sjostrand et al.
+[2006].
+It incorporates nuclear parton distribution functions (PDFs) from LHAPDF5 Whalley et al.
+[2005], intranuclear scattering via DPMJet Roesler et al.
+[2000], and the geometric density of nucleons through PyQM Dupré [2011].
+Nuclear remnant de-excitations and decays are processed using FLUKA Ferrari et al.
+[2005].
 We determined the correlation functions from these simulations by applying Eq.
@@
 The value of $C_{0}$ used for normalizing the simulated correlation functions was determined using the deuterium simulations and is independent of the value used for data.
+However, they are close to one another: $C_{0}^{\rm data}\approx 7.5$ and $C_{0}^{\rm BeAGLE}\approx 8.1$, indicating that the prediction for the average number of ejected protons per leading-pion event in deuterium is within 10% of that of the data.
 We find that the trends are similar to those of the data; however, there are numerical discrepancies.
@@
 The horizontal caps in the error bars represent the systematic uncertainties, while the outer bars represent the total systematic and statistical uncertainty (added in quadrature).
-The results are compared to the predictions from the BeAGLE Chang:2022hkt, eHIJING PhysRevD.110.034001, and GiBUU Buss:2011mx models, as well as analogous measurements of the dipion channel from Ref.
-PhysRevC.111.035201 (gray, open squares).
+The results are compared to the predictions from the BeAGLE Chang et al.
+[2022], eHIJING Ke et al.
+[2024], and GiBUU Buss et al.
+[2012] models, as well as analogous measurements of the dipion channel from Ref.
+Paul and others [2025] (gray, open squares).
 The broadenings for the dipion channel were reported in Ref.
-PhysRevC.111.035201 with respect to deuterium, and have been re-calculated with respect to carbon in order to compare them to the results of this work.
+Paul and others [2025] with respect to deuterium, and have been re-calculated with respect to carbon in order to compare them to the results of this work.
 From the correlation functions, we determined the RMS widths, $\sigma$, integrated in $\Delta Y$, which we show in the top left panel of Fig.
@@
 We note that $\sigma$ increases monotonically with respect to increasing nuclear size, with values of about 1.13, 1.32, 1.41, and 1.45 for D, C, Fe, and Pb, respectively.
-This increase is further reflected in the azimuthal broadenings, $b$, which increase as a function of AA, as shown in the lower left panel of the same figure.
-This trend is similar to what was observed in the dipion measurements PhysRevC.111.035201 (gray open squares in Fig.
+This increase is further reflected in the azimuthal broadenings, $b$, which increase as a function of $A$, as shown in the lower left panel of the same figure.
+This trend is similar to what was observed in the dipion measurements Paul and others [2025] (gray open squares in Fig.
 3).
 We note that this increase in width with respect to nuclear size is steeper in this work than in Ref.
-PhysRevC.111.035201.
+Paul and others [2025].
 This seems to suggest that low-energy protons undergo stronger final-state interactions than secondary pions.
-In the context of some models, such as the GiBUU model Buss:2011mx, when secondary pions are produced, they have a “pre-hadronic” phase when they are color neutral but are not fully formed.
+In the context of some models, such as the GiBUU model Buss et al.
+[2012], when secondary pions are produced, they have a “pre-hadronic” phase when they are color neutral but are not fully formed.
 During this phase, the “pre-hadrons” have a weaker interaction with the nuclear material than after they have fully formed.
@@
 The widths increase monotonically as a function of $\Delta Y$ (which was also observed for the dipion channel in Ref.
-PhysRevC.111.035201, however we observe that the slope of this dependence is larger for the $\pi p$ channel than for the dipion channel).
+Paul and others [2025], however we observe that the slope of this dependence is larger for the $\pi p$ channel than for the dipion channel).
 The broadening is consistently negative (up to $\approx$-0.8) for deuterium and positive for iron (up to $\approx 0.6$) and lead (up to $\approx 0.7$), with the exception of one bin in which the uncertainties are very large.
@@
 However, the agreement with the data is not as good within slices in $\Delta Y$ as in the integrated results.
-We also compared our results for the widths and broadenings to the predictions of the eHIJING model PhysRevD.110.034001 (version 1.0) and the GiBUU model Buss:2011mx (2021 release) in Fig.
+This appears to be related to the discrepancy mentioned above, where the predicted correlation functions appear to be shifted in $\Delta Y$ relative to the data.
+We also compared our results for the widths and broadenings to the predictions of the eHIJING model Ke et al.
+[2024] (version 1.0) and the GiBUU model Buss et al.
+[2012] (2021 release) in Fig.
 3.
-We also provide a comparison between these models’ predictions for the correlation functions and the ratio, $R$, in the Supplementary Materials supplementary.
-In eHIJING, the hard-scattering process, along with initial- and final-state fragmentation, is simulated using PYTHIA8 bierlich2022comprehensive.
+We also provide a comparison between these models’ predictions for the correlation functions and the ratio, $R$, in the Supplementary Materials 1.
+In eHIJING, the hard-scattering process, along with initial- and final-state fragmentation, is simulated using PYTHIA8 Bierlich et al.
+[2022].
 The model incorporates nuclear modifications to PDFs, fragmentation, and parton showering.
@@
 The GiBUU model incorporates treatment of final-state interactions, absorption, and production mechanisms with elastic and inelastic channels.
-The GiBUU model describes the single-hadron and di-hadron data reasonably well from CLAS CLAS:2021jhm, CLAS:2022asf, PhysRevC.111.035201 and HERMES HERMES:2003icw, Airapetian:2007vu, Airapetian:2011jp.
+The GiBUU model describes the single-hadron and di-hadron data reasonably well from CLAS Moran and others [2022], Paul and others [2022, 2025] and HERMES Airapetian and others [2003, 2007, 2011].
 We modified the default values of several parameters related to parton distributions and fragmentation in free nucleons, based on a tuning study that used an independent proton-target dataset from Ref.
-CLAS:2022sqt, and which was further validated with dipion correlations in Ref.
-PhysRevC.111.035201.
+Avakian and others [2023], and which was further validated with dipion correlations in Ref.
+Paul and others [2025].
 This tuning does not modify the nuclear effects.
@@
 The azimuthal‑peak width grows systematically with nuclear size, and the proton yield per pion also rises but it appears to saturate for the heaviest targets.
-These patterns are consistent with a picture in which hadron absorption and interactions in the hadronic cascade widens the correlation and the larger pool of protons in heavier nuclei enhances the yield.
-Compared with earlier dipion measurements PhysRevC.111.035201, the $\pi p$ channel shows a markedly stronger dependence on both nuclear size and rapidity separation.
-The trends are qualitatively reproduced by the state-of-the-art eAeA event-generator models BeAGLE, eHIJING, and GiBUU, but quantitative differences remain, highlighting opportunities to refine the treatment of nuclear fragmentation.
-Our results establish a benchmark for forthcoming di-hadron and slow-nucleon production studies in DIS, as well as for the development of corresponding event-generator models for experiments at Jefferson Lab Burkert:2018nvj, Burkert:2020akg, Arrington:2021alx and at future Electron–Ion Colliders Accardi:2012qut, Anderle:2021wcy.
+These patterns are consistent with a picture in which hadron absorption and interactions in the hadronic cascade widen the correlation and the larger pool of protons in heavier nuclei enhances the yield.
+Compared with earlier dipion measurements Paul and others [2025], the $\pi p$ channel shows a markedly stronger dependence on both nuclear size and rapidity separation.
+The trends are qualitatively reproduced by the state-of-the-art $eA$ event-generator models BeAGLE, eHIJING, and GiBUU, but quantitative differences remain, highlighting opportunities to refine the treatment of nuclear fragmentation.
+Our results establish a benchmark for forthcoming di-hadron and slow-nucleon production studies in DIS, as well as for the development of corresponding event-generator models for experiments at Jefferson Lab Burkert [2018], Burkert and others [2020], Arrington and others [2022] and at future Electron–Ion Colliders Accardi and others [2016], Anderle and others [2021].
 7 Code and Data Availability
 We provide in Ref.
-zenodoRivet the code we used to extract the $(\Delta\phi,\Delta Y)$ histograms from the output files of event generators (in order to produce the correlation functions) in the form of a module in the Rivet framework 10.21468/SciPostPhys.8.2.026.
-The values of the correlation functions, the ratio $R$, RMS widths, and broadenings reported in this work can be found in the Supplementary Materials supplementary.
+Paul the code we used to extract the $(\Delta\phi,\Delta Y)$ histograms from the output files of event generators (in order to produce the correlation functions) in the form of a module in the Rivet framework Bierlich et al.
+[2020].
+The values of the correlation functions, the ratio $R$, RMS widths, and broadenings reported in this work can be found in the Supplementary Materials 1.
 8 Acknowledgements
@@
 We thank Kai Gallmeister for help in setting up the GiBUU event generator and Weiyao Ke for help in setting up the eHIJING event generator.
-This work is supported by the Chilean Agencia Nacional de Investigacion y Desarollo (ANID), FONDECYT Grants No.
+This work is supported by the Chilean Agencia Nacional de Investigación y Desarrollo (ANID), FONDECYT Grants No.
 1221827 and No.